an elephant of mass 500kg stands with all four legs on the ground esch foot has area of 0.08m^2.calculate the pressure exerted by the elephant on the ground(g=20m/s^2)


1
Expert's answer
2022-03-03T12:05:13-0500

The pressure

"P=\frac{F}{A}=\frac{mg}{A}"

"P=\frac{500*10}{4*0.08}=15625\:\rm Pa"